OpenStudy (anonymous):

if h(x)=(fog)(x) and h(x)√4:(x+7) FIND g(x) if f(x)=√4:(x+1).

OpenStudy (anonymous):

what does \[\sqrt{4:(x+1)}\] mean?

OpenStudy (anonymous):

if it is \[h(x)=\sqrt{4(x+7)}\] and \[f(x)=\sqrt{4(x+1)}\] then \[g(x)=x+6\]

OpenStudy (anonymous):

\[\sqrt[4]{x+7}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

sorry :/

OpenStudy (anonymous):

so \[h(x)=\sqrt[4]{x+7}\] and \[f(x)=\sqrt[4]{x+1}\] then \[g(x)=x+6\]

OpenStudy (anonymous):

thank you!